# Table Foundation Models: on knowledge pre-training for tabular learning

### v1

@@ 2 The unfolding of tabular foundation models
Compared to neural networks, tree-based models have a lead start for tabular learning: their inductive biases match well properties of tabular data [Grinsztajn et al., 2022]. Progress in dedicated neural architectures has recently been closing the gap for large-enough datasets [Borisov et al., 2022, Ye et al., 2024]. But tabular foundation models bring the promise of benefits for data of small to moderate size. van Breugel and van der Schaar [2024] argue they should be a research priority, calling for developing properties important to tabular applications such as cross-dataset modeling as well as handling tables with different types, e.g., numbers and strings. Recent progress on this agenda has required overcoming many table-specific challenges.
@@ 3.2 Knowledge pre-training from large knowledge bases
The transformer architecture of TARTE is specified as follows: We set $3$ self-attention layers with $24$ multi-head attentions, $768$ hidden dimension, and 2 04820482\,0482 048 feed-forward dimension per layer. For the projection layers for contrastive learning, we set two linear layers with hidden and output dimensions of 2 04820482\,0482 048 and $768$, respectively. The resulting model contains over $25$ million trainable parameters. The batch size is set as $512$ in which $256$ entities are randomly selected with one additional positive for each. The total number of steps for training is 200 000200000200\,000200 000, with the AdamW optimizer and the cosine scheduler. The learning rates were set as $lr_{min}=10^{-8}$, $lr_{max}=10^{-6}$ of warm-up over the first 2 00020002\,0002 000 steps, followed by a linear decay in learning rate schedule. The probability for all dropout layers was set as $0.1$.
@@ 3.3 Learning with the backbone: fine-tuned, or frozen, combined with another model
Similar to the sentence-transformers [Reimers and Gurevych, 2019] on LLMs, TARTE can be used to generate meaningful embeddings for table entries. From the preprocessed input of a downstream table, the embedded representation of the readout token, $T$, can be used with any machine learning model as a pipeline to make predictions on unseen data.
@@
As TARTE is pre-trained from large knowledge bases, the embeddings of the TARTE featurizer potentially hold implicit background information. Such prior knowledge can be useful, especially in few-shot settings; but when the original table provides sufficient information for learning, the background information, on its own, becomes less useful for inference (see Figure 3). Yet, we argue that the implicit prior information continues to be useful when combined appropriately with a complementary tabular model. To accompany the prior to tabular models, we formulate a boosting strategy: the base model for tabular data is boosted with a model that fits the (train) residuals of the base model with TARTE embedded features. For efficiency, we use TARTE with a simple Ridge regression.
@@ 3.4 Differences to the CARTE approach
Third, TARTE pre-training reflects better downstream tasks by handling different data types and better pre-training data sources (detailed above), carefully preprocessing the batch samples, such as trimming to match the number of columns and limiting redundant relations. In addition, TARTE carefully controls the learning rates and the number of training steps to avoid overfitting the model to the pre-train data.
@@ 4.1 Experimental set up: tabular learning
Datasets
We use the benchmark from Kim et al. [2024]: 40 regression and 11 classification datasets [^3] Datasets available at https://huggingface.co/datasets/inria-soda/carte-benchmark. These tables come with informative columns and discrete entries.
Methods
We evaluate post-training paradigms of TARTE (subsection 3.3) and the best performers in related benchmarks: the leading gradient-boosted tree models, XGBoost [XGB, Chen and Guestrin, 2016] and CatBoost [Prokhorenkova et al., 2018]; pre-trained models, TabPFNv2 [Hollmann et al., 2025] and CARTE [Kim et al., 2024]; simple linear model Ridge from scikit-learn [Pedregosa et al., 2011].
In terms of data preparation, most of the compared models provide native handling of diverse data types. However, for models that explicitly require numerical tables, we rely on heuristics provided by the TableVectorizer (TabVec) functionality of the skrub software. While TabPFNv2 handles readily the input tables, we optionally combine it with the TableVectorizer, for better handling of high-cardinality string and datetime columns.
@@ Methods
  • TARTE – : TARTE as table featurizer with Ridge, XGB, and TabPFNv2 as prediction models.
@@ 4.2 TARTE boosts learning, and can be computationally efficient
Figure 3 (top) shows learning curves of normalized prediction scores as a function of sample sizes. The scores are normalized across all train sizes per dataset, with $1$ as the best and 00 as the worst performing model. Pre-trained models (TARTE, CARTE, and TabPFNv2) generally perform better, with variants of TARTE as the best performing models, regardless of the sample sizes. Critical difference diagrams (Figure 3, middle) show that the benefits brought by TARTE are significant [^4] Conover post hoc test after a Friedman test for pairwise significance [Conover, 1999]..
@@
TARTE–B consistently improves the base model for both TabPFNv2 and XGB. This suggests that TARTE embeddings complement base models, bringing prior information acquired from knowledge pre-training. In addition, TARTE featurizer on its own gives good table preparation. Compared to TabVec (dashed lines), TARTE featurizer can markedly increases performance, bringing, for instance, ridge to a very good position.
@@ 4.3 Good knowledge pre-training, adapted to downstream tasks, is important
TARTE embeddedings boost prediction. But what drives this boost? Is it the inductive bias of the architecture, or is the pre-training on the knowledge base actually important? To answer this question, we investigate TARTE pre-training on small tables ($n=32$ to 1 02410241\,0241 024), where the effect of TARTE is most eminent. Figure 6 presents the critical difference diagram for Ridge fitted with embeddings from different pre-training schemes. Here, “Random” corresponds to random weights (no pre-training), and “MinHash” replaces FastText by MinHash encoding, from the skrub package, that captures morphological but not semantic similarities of strings [Cerda and Varoquaux, 2020].
Figure 6: Comparison of different pre-training schemes. A ridge is fitted with embeddings from different schemes. “Random” stands for random weights (no pre-training), and “MinHash” corresponds to the skrub’s MinHash encoder in replacement of FastText. TARTE is pre-trained with different knowledge bases: YAGO3, and our enriched YAGO4.5 (TARTE–Enriched YAGO4.5 is called TARTE–Ridge elsewhere).
First, comparing MinHash to TARTE reveals the importance of FastText. A simple use of FastText with non pre-trained weights, TARTE–Random (or CARTE–Random), already exhibit relatively strong performances. As FastText captures semantic similarities, its combination with a suitable transformer architecture (subsection 3.1 or CARTE) forms an inductive bias of smoothness: similar tables have similar representations.
Going further, a proper pre-training of TARTE improves performances (TARTE–Enriched YAGO4.5). For CARTE, pre-training actually hurts performance without fine-tuning (CARTE–Random to CARTE-YAGO3): these representations lack out-of-the-box re-usability. On the other hand, TARTE improves with knowledge pre-training, and all the more with more relevant pre-training data (moving from YAGO3 to Enriched YAGO4.5, and preprocesssing to better reflect the downstream tables, see subsection 3.2).
Table 1: Factors of success for TARTE. Coefficients, and corresponding confidence intervals, of a linear model explaining the improvement of TARTE–Ridge over TabVec–Ridge across datasets (R-squared: 0.22). Negative values mean that greater values for the corresponding meta-feature are associated to less benefits of TARTE, for instance longer strings.
@@
We find that TARTE works less well on tables with long string entries (likely a limitation of FastText embedings). However, it brings more benefits for tables with many strings considered as inliers to the pre-training source (using a One-Class SVM fitted on the embeddings of the pre-training strings to define an inlier score). It also works better with fewer low-cardinality columns (trees shine on these), with many datetime columns, with strings similar across rows, with more numerical or high-cardinality columns, and with more data. Overall, we find that TARTE brings benefits on more complex tables, with many complex strings, but particularly so if these look like strings seen during pre-training.
@@ 4.4 Specializing to a domain: fine-tuned TARTE gives representations that transfer
We investigate whether reusing TARTE representations from fine-tuned models facilitates subsequent learning. We evaluate TARTE–Ridge and TARTE–B with TabPFNv2 and XGB. Here, TARTE–Ridge denotes a Ridge model fitted on fine-tuned TARTE embeddings and TARTE–B sequentially fits the residuals with domain specialized representations (see subsection 3.3). As baselines, we consider top-performing tabular models that provide native handling of missing values. Indeed, even with differing columns across tables, these models can be fitted with a stacked table. We also include CARTE multi-table (CARTE-MT), the only baseline, to our knowledge, that learns across tables without correspondences [Kim et al., 2024]. Note that the setup for TARTE and baseline models is slightly different. Baselines fit all tables, sources and target, jointly; while the TARTE model does not need access to the source tables to be reused for transfer. We consider each domain one after the other, and for each, we consider every combination of source tables and a target table: source tables to specialize the model, and target table to evaluate the corresponding learner. See appendix B.3 for details.
@@
TARTE stays efficient with multiple source tables
We evaluate every possible source–target combinations with TARTE–B–TabPFNv2(DS) and CARTE-MT. Figure 8 gives the improvements of respective models compared to the best performing model on single tables (TARTE–B–TabPFNv2), with respect to the runtime. We see that both models benefit from source tables, with larger gains for TARTE–B–TabPFNv2(DS). Compared to [Kim et al., 2024], the baseline here is much more performant, and thus it is more difficult to improve of top of it. Concerning the runtime, CARTE-MT requires pairwise fitting for each source, markedly increasing compute costs even for small increase in train-size. TARTE, however, is more efficient. Similar to the case of single source, the runtime of TARTE is primarily dominated by the training time of source tables (20 mn). Therefore, reusing an existing domain-specialized models imposes far less compute cost, putting TARTE in a better position for transfer within the specific domain.
@@ A.1 Modeling with column–cell pairs: context-aware for transformers
A.2 Pre-train data
@@
A.3 Pre-training procedures
@@ B.1 Training procedures
For evaluation on single tables, the train-size for each tables varied from $32$, $64$, $128$, $256$, $512$, 1 02410241\,0241 024, and 10 0001000010\,00010 000; remaining data was set as the test set. The performance was measured with $R^{2}$ score for regression and the Area Under Receiver Operating Curve (AUROC) for classification tasks. For the runtime, it measures the total time for data preparation, hyperparameter optimization, and prediction. Overall, the results were recorded on $10$ different train/test splits for each dataset.
B.2 Details on multivariate analysis
B.3 Experimental set-up for domain specialization
Datasets
@@ D.1 Results on small tables
Figure 9 presents the results for all train-sizes in few-shot setting. We include additional baselines of Random Forest (RF), the classical Multilayer Perceptron (MLP), and TARTE–CatBoost using the default hyperparameters. TARTE, through various post-training schemes, outperforms the baselines regardless of train-sizes. In particular, TARTE featurizer outperforms TableVectorizer (except for TabPFNv2) for a given learner in small training samples ($n\leq 256$). For TARTE–TabPFNv2, it underperforms likely due to that the dimension of TARTE embeddings was $256$ for TabPFNv2 and the resulting data representations do not have a distribution matching to that of synthetic tables from TabPFNv2 pre-training.

### v2

@@ 2 The unfolding of tabular foundation models
Compared to neural networks, tree-based models have a lead start for tabular learning: their inductive biases match well the properties of tabular data [Grinsztajn et al., 2022]. Progress in dedicated neural architectures has recently been closing the gap for large-enough datasets [Borisov et al., 2022, Ye et al., 2024]. But tabular foundation models bring the promise of benefits for data of small to moderate size. van Breugel and van der Schaar [2024] argue they should be a research priority, calling for developing properties important to tabular applications such as cross-dataset modeling as well as handling tables with different types, e.g., numbers and strings. Recent progress on this agenda has required overcoming many table-specific challenges.
@@ 3.2 Knowledge pre-training from large knowledge bases
The transformer architecture of TARTE is specified as follows: We set three self-attention layers with $24$ multi-head attentions, $768$ hidden dimension, and 2 04820482\,0482 048 feed-forward dimension per layer. For the projection layers for contrastive learning, we set two linear layers with hidden and output dimensions of 2 04820482\,0482 048 and $768$, respectively. The resulting model contains over $25$ million trainable parameters.
@@ 3.3 Learning with the backbone: fine-tuned, or frozen, combined with another model
Similar to the sentence-transformers [Reimers and Gurevych, 2019] on LLMs, TARTE can be used to generate meaningful embeddings for table entries. The preprocessed input of a downstream table is passed through the frozen backbone of TARTE, and the embedded representation of the readout element $T$ (similar to the [CLS] token in LLMs) can be used with any machine learning model as a pipeline to make predictions on unseen data.
@@
As TARTE is pre-trained from large knowledge bases, the embeddings of the TARTE featurizer potentially hold implicit background information. Such prior knowledge can be useful, especially in few-shot settings; but when the original table provides sufficient information for learning, the background information, on its own, becomes less useful for inference (see Figure 3). Yet, we argue that the implicit prior information continues to be useful when combined appropriately with a complementary tabular model. To accompany the prior to tabular models, we formulate a boosting strategy: the base tabular model with the original table is ensembled with a model that fits the (train) residuals of the base model with TARTE embedded features. For efficiency, we use TARTE with a Ridge regression.
@@ 3.4 Differences to the CARTE approach
Third, TARTE reflects better downstream tasks by handling different data types and better pre-training data sources (see subsection 3.2 and appendix B.1), carefully preprocessing the batch samples, such as trimming to match the number of columns and limiting redundant relations. In addition, TARTE carefully controls the training procedures (see appendix A.4).
@@ 4.1 Experimental set up: tabular learning
Datasets and methods
We use the benchmark from Kim et al. [2024]: 40 regression and 11 classification datasets [^3] Datasets available at https://huggingface.co/datasets/inria-soda/carte-benchmark. These tables come with informative columns and discrete entries. For comparing methods, evaluate post-training paradigms of TARTE (subsection 3.3) and the best performers in related benchmarks: the leading gradient-boosted tree models, XGBoost [XGB, Chen and Guestrin, 2016] and CatBoost [Prokhorenkova et al., 2018]; neural-network model, RealMLP [Holzmüller et al., 2024]; pre-trained models, TabPFNv2 [Hollmann et al., 2025] and CARTE [Kim et al., 2024] with fine-tuning and boosting (CARTE–B–); simple linear model Ridge from scikit-learn [Pedregosa et al., 2011].
In terms of data preparation, most models provide native handling of diverse data types. However, for models that explicitly require numerical tables, we rely on heuristics provided by the TableVectorizer (TabVec) functionality of the skrub software. While TabPFNv2 handles readily the input tables, we optionally combine it with the TableVectorizer, for better handling of high-cardinality string and datetime columns.
@@
  • TARTE – : TARTE as table featurizer. We consider Ridge, XGB, and TabPFNv2 as prediction models.
@@ 4.2 TARTE boosts learning, and can be computationally efficient
Figure 3 (top) shows learning curves of normalized prediction scores as a function of sample sizes. The scores are normalized across all train sizes per dataset, with $1$ as the best and 00 as the worst performing model. Pre-trained models (TARTE, CARTE, and TabPFNv2) generally perform better, with variants of TARTE as the best performing models, regardless of the sample sizes. Critical difference diagrams [^4] Critical difference diagrams display average rank across models with crossbars depicting the two having no statistically significant difference based on Conover post hoc test after a Friedman test for pairwise significance [Conover, 1999]. (Figure 3, middle) show that the benefits brought by TARTE are significant.
@@
In general, boosting a base model with pre-trained embeddings (TARTE–B and CARTE–B) improves prediction. However, as TARTE is better pre-trained (see subsection 4.3), its embeddings complement better the base models, bringing prior information acquired from knowledge pre-training. The best performances require a strong base tabular model, as with TabPFNv2 and XGB. On the other hand, TARTE–FT does not have such requirements, but it comes with computation burdens, with extensive hyperparameter tuning [^5] Due to the computation costs, the search space for TARTE-FT is limited (see Table 3). Finally, TARTE featurizer is a good table preparator: compared to TabVec (dashed lines), it can markedly increase performance, bringing, for instance, ridge to a good position.
@@ 4.3 Good knowledge pre-training, adapted to downstream tasks, is important
TARTE embeddedings boost prediction. But what drives this boost? Is it the inductive bias of the architecture, or is the pre-training on the knowledge base actually important? To answer this question, we investigate TARTE pre-training on small tables ($n=32$ to 1 02410241\,0241 024), where the effect of TARTE is most eminent. Figure 6 presents the critical difference diagram for Ridge fitted with embeddings from different schemes. Here, CARTE with “New sampling” corresponds to CARTE architecture with TARTE pre-training schemes, and “MinHash” replaces FastText by MinHash encoding, that only captures morphological similarities of strings [Cerda and Varoquaux, 2020].
Figure 6: Ablating architecture, pre-training, and preprocessing components. A ridge is fitted with embeddings from different schemes: random weights (no pre-training); replacement of FastText with skrub’s “MinHash” encoder; TARTE and CARTE pre-trained with YAGO3 and Enriched YAGO4.5; CARTE with TARTE pre-training schemes (New sampling); TARTE without datetime detection or column information.
First, comparing MinHash to TARTE reveals the importance of FastText. A simple use of FastText with random (non pre-trained) weights already exhibit relatively strong performances. As FastText captures semantic similarities, its combination with a suitable transformer architecture (subsection 3.1 or CARTE) forms an inductive bias of smoothness: similar tables have similar representations.
Going further, a proper combination of pre-train data and procedures, the architecture, and preprocessing of tables improves knowledge pre-training (TARTE with Enriched YAGO4.5). For CARTE representations used without fine-tuning, pre-trained representations perform less well compared to random weights, regardless of the pre-train data: these representations lack out-of-the-box re-usability. TARTE pre-training schemes on CARTE (New sampling) helps but the performance does not match that of TARTE, highlighting the various sources of improvements through TARTE (see subsection 3.2). In addition, without datetime detection or column information, properly pre-trained weights still provide competitive representations. This suggests that TARTE can perform well even on tables without meaningful column names.
Table 1: Factors of success for TARTE. Coefficients and their confidence intervals of a linear model explaining the improvement of TARTE–Ridge over TabVec–Ridge across datasets (R-squared: 0.22). Negative values mean that greater values for the corresponding meta-feature are associated to less benefits of TARTE. Performance decrease for long strings suggest a limitation of the language model, FasText. Importance of inlier probability shows the need for pretraining to cover well downstream terminology.
@@
We find that TARTE works less well on tables with long string entries (likely a limitation of FastText embedings, as discussed in appendix A.2). However, it brings more benefits for tables with many strings considered as inliers to the pre-training source (using a One-Class SVM fitted on the embeddings of the pre-training strings to define an inlier score). TARTE also works better with fewer low-cardinality columns (trees shine on these), with many datetime columns, with strings similar across rows, with more numerical or high-cardinality columns, and with more data. Overall, we find that TARTE brings benefits on more complex tables, with many complex strings, but particularly so if these look like strings seen during pre-training.
@@ 4.4 Specializing to a domain: fine-tuned TARTE gives representations that transfer
We investigate whether reusing TARTE representations from fine-tuned models facilitates subsequent learning. We evaluate TARTE–Ridge and TARTE–B with TabPFNv2 and XGB. Here, TARTE–Ridge denotes a Ridge model fitted on fine-tuned TARTE embeddings and TARTE–B sequentially fits the residuals with domain specialized representations (see subsection 3.3). As baselines, we consider top-performing tabular models that provide native handling of missing values. Indeed, even with differing columns across tables, these models can be fitted with a stacked table. We also include CARTE multi-table (CARTE-MT), the only baseline, to our knowledge, that learns across tables without correspondences [Kim et al., 2024]. Note that the setup for TARTE and baseline models is slightly different. Baselines fit all tables, sources and target, jointly; while the TARTE model does not need access to the source tables to be reused for transfer. We consider each domain one after the other, and for each, we consider every combination of source tables and a target table: source tables to specialize the model, and target table to evaluate the corresponding learner. As datasets, we select from the previous set of tables group of tables within the same domain, but acquired in different settings [Kim et al., 2024]. See appendix B.3 for details.
@@
TARTE stays efficient with multiple sources
We evaluate every possible source–target combinations with TARTE–B–TabPFNv2(DS) and CARTE-MT. Figure 8 gives the improvements of respective models compared to the best performing model on single tables (TARTE–B–TabPFNv2), with respect to the runtime. Both models benefit from source tables, with larger gains for TARTE–B–TabPFNv2(DS). Compared to [Kim et al., 2024], the baseline here is much more performant, and thus it is more difficult to improve. Concerning the runtime, CARTE-MT requires pairwise fitting for each source, markedly increasing compute costs even for small increase in train-size. TARTE, however, is more efficient: the runtime of TARTE is dominated by the training time of source tables (20 mn). Therefore, reusing domain-specialized models imposes far less compute cost, putting TARTE in a better position for transfer within the specific domain.
@@ A.1 Modeling with column–cell pairs: context-aware for transformers
A.2 Language model used in TARTE
We use FastText embeddings [Mikolov et al., 2017], which hold some limitations for long string entries (subsection 4.3). Yet, string (text) entries in tables mostly contain only a few words: across 51 datasets, the median ratio of unique entries with more than 10 words is 0.025. An interesting alley to explore more sophisticated language models, possibly using different language models depending on different structure of strings (e.g., simple words or sentences).
A.3 Pre-train data
@@
A.4 Pre-training procedures
Training specifications
The batch size is set as $512$ in which $256$ entities are randomly selected with one additional positive for each. The total number of steps for training is 200 000200000200\,000200 000, with the AdamW optimizer and the cosine scheduler. The learning rates were set as $lr_{min}=10^{-8}$, $lr_{max}=10^{-6}$ of warm-up over the first 2 00020002\,0002 000 steps, followed by a linear decay in learning rate schedule. The probability for all dropout layers was set as $0.1$.
@@ B.1 Training procedures
For evaluation on single tables, the train-size for each tables varied from $32$, $64$, $128$, $256$, $512$, 1 02410241\,0241 024, and 10 0001000010\,00010 000; remaining data was set as the test set. Out of 51 datasets, 27 datasets contain sufficient data points for the train-size of 10 0001000010\,00010 000. The performance was measured with $R^{2}$ score for regression and the Area Under Receiver Operating Curve (AUROC) for classification tasks. For the runtime, it measures the total time for data preparation, hyperparameter optimization, and prediction. Overall, the results were recorded on $10$ different train/test splits for each dataset.
@@
B.2 Details on multivariate analysis
B.3 Experimental set-up for domain specialization
Datasets
@@ D.1 Results on small tables
Figure 9 presents the results for all train-sizes in few-shot setting. We include additional baselines of Random Forest (RF), the classical Multilayer Perceptron (MLP), and TARTE–CatBoost using the default hyperparameters. TARTE, through various post-training schemes, outperforms the baselines regardless of train-sizes. In addition, TARTE featurizer outperforms TableVectorizer (except for TabPFNv2) for a given learner in small training samples ($n\leq 256$). For TARTE–TabPFNv2, it underperforms likely due to that the dimension of TARTE embeddings was $256$ for TabPFNv2 and the resulting data representations do not have a distribution matching to that of synthetic tables from TabPFNv2 pre-training.